OpenStudy (anonymous):

The Martians ask you to explain one last thing, Ultimate Math Ambassador. Assign any number to x. Using complete sentences, explain whether f(g(x)) and g(f(x)) will always result in the same number. You will use the inverse function that you created in problem number 5 for g(x).

OpenStudy (ddcamp):

From problem 5, what are f(x) and g(x)?

OpenStudy (anonymous):

(x-10)/2=g(x)

OpenStudy (anonymous):

F(x)=2x+10

OpenStudy (anonymous):

do i plug in my value for g(x) into my f(x)?

OpenStudy (anonymous):

g(x)= (x-10) 2 g(20)= (20-10) 2 g(20)= 10/2 g(20)=5 i think i did that correctly..

OpenStudy (ddcamp):

Yes, I think that would work. The other thing you could do is replace each x in f(x) with g(x): \[2x+10 \rightarrow 2(\frac{ x-10 }{ 2 }) + 10\]

OpenStudy (ddcamp):

Just to show that f(g(x)) = x and g(f(x)) = x in all situations.

OpenStudy (anonymous):

so do i plug in my x value (20) into that equation?

OpenStudy (anonymous):

never mind thank you :)